# Compute Optimal Scaling of Skills: Knowledge vs Reasoning

### v2

@@ Results
In Figure Figures 1(a) and 1(b), respectively, we show average results for the held-out code and knowledge QA datasets (in red). For comparison, we overlay the APE CO curve reported by Dubey et al. (2024) (in black). The figure shows a clear difference between knowledge QA and code: while knowledge QA prefers capacity compared to the APE curve – especially at lower compute scales – code tends to be more data-hungry. This pattern is present across all datasets belonging to the hypothesis split on which we first observed this difference, as well as for the held-out split. We show individual plots for all datasets in Figures 11, 12, 13 and 14.
To make the pattern more explicit across different scales, we consider the distribution of parameter count ‘residuals,’ or log-scale differences between the APE optima and the skill-optima for each skill. If at compute scale $B$, the residual is larger than 0, we say that the skill is capacity-hungry at $B$; if it is smaller than 0, we call it data-hungry at $B$. The resulting distributions, shown in Figure 1(c), confirm the pattern we observed before: the parameter count residuals for code are shifted substantially to the left with respect to the parameter count residuals for knowledge QA, indicating that the latter skill is more capacity-hungry.
@@
In Figure 2, we show the results for the code and knowledge held-out datasets. Unsurprisingly, increasing the amount of skill-dependent data (lighter colors in the plot), improves performance for that specific skill. [^3] Increasing code data improved losses on code datasets beyond the next compute scale on the canonical datamix. This pattern is much clearer for code than for knowledge. We hypothesise that this is because code datapoints are easily identified and contain only code, whereas knowledge datasets are usually more dispersed as well as more difficult to label automatically. More relevant to our research questions is the fact that not only the losses, but also the COs shift. Again specifically for code, this pattern is clear in both the hypothesis and held-out sets – the more code data is included, the more capacity-hungry the CO becomes. For reference, in Appendix E, we provide individual plots for all code and knowledge QA datasets of all splits, across all ablations.
@@
Finally, we use the same data to investigate the data-hunger of knowledge QA and code independently of the proportion of skill relevant data. To do so, we plot the CO parameter counts for each skill as a function of the proportion of skill relevant data. In Figure 3, we show this for both knowledge QA and code in the same plot. [^4] The axis is aligned on proportions of skill-relevant data, not on the datamixes themselves. For knowledge QA datasets (purple lines), an x-value of 0.4 corresponds to a datamix with 40% knowledge, whereas for code datasets (green lines), it corresponds to results for the datamix with 40% code. From this figure it is clear that even correcting for the proportion of skill relevant data, knowledge is substantially more capacity-hungry than code, and increases its capacity-hunger more quickly as the proportion of skill relevant data increases. We hypothesise that this phenomenon, which demonstrates a fundamental difference between knowledge and code, is related to the fact that knowledge is harder to compress than code, requiring more capacity to memorise facts. In line with this hypothesis, we observe a larger difference between knowledge and code when only Python code datasets are considered, as can be seen in Figure 4. We hypothesise that models at this scale may unable to succesfully compress ‘low-resource’ programming languages, though further research is required to confirm this.555 We also see in Figures 2 and 1 and Appendices E and I that datasets quantifying code have a lower loss value than those quantifying knowledge QA– across the board – despite large proportions of knowledge, which could suggest that code has lower unmodelable entropy compared to knowledge.
@@ 4.3.1 Aligning skill-COs
In Figure 5(b), we show the NLL for all validation sets, using the same axis as in Figure 5(a). The vertical dashed line indicates the optimal point from Figure 5(a). A first striking observation is that all the chosen validation sets track either knowledge QA or code quite closely: while the loss on FineWeb and FineWeb-Edu monotonically increases with the amount of code data, the opposite is true for The Stack. We do not see evidence of a mix of knowledge QA and code in the validation sets, which might have presented itself as a U-shaped curve, rather than the monotonically increasing or decreasing curves that we observe. To compute appropriate scaling laws, none of these validation sets would thus have resulted in COs aligned with both skills.
@@ 4.3.2 The impact of the validation set
Lastly, we study more explicitly how much variation in the predicted COs is to be expected depending on the choice of validation set. In Figure 6(a), considering the The stack validation set as ‘baseline,’ we can see that for some datamixes, the optimal parameter count is off by almost 50%. In Figure 6(c), we see that, for the smallest compute scale, for the canonical datamix there is over a 30% difference in optimal parameter count between the different validation sets. While the relative difference is not as stark at higher compute scales, it still exceeds 10% for the 3 largest compute scales.
With that, our experiments show that choosing a validation set that adequately represents a mix of what the final model should capture – or perhaps even a validation set that more directly measures those skills, such as in our work – is critical to finding the right COs. While the difference is more pronounced at smaller scales, it persists even on the largest compute scale in our experiments. On the other hand, Dubey et al. (2024) points out that IsoFLOP curves get flatter at high compute scales, so the tradeoff between model capacity and dataset size might not matter as much. In one of their other results, Dubey et al. (2024) uses a fitted sigmoid of the loss to estimate accuracies. If one were to take their analysis a step further by applying a sigmoid to an entire IsoFLOP curve to produce an IsoFLOP curve of accuracy rather than NLL, we hypothesise that this could make curves appear less flat if low compute scales are in the flat region of the sigmoid and high compute scales are in the linear region. We leave this intriguing question to future work.
@@ 5 Related work
The first large-scale empirical study of statistical scaling laws of neural networks was done by Hestness et al. (2017), who studied scaling laws for a variety of tasks using the LSTM Hochreiter and Schmidhuber (1997) and CNN Lecun et al. (1998) architectures. Their paper used power law fits to their experimental data, and they found that the only way to materially change the power law exponent was to change the task, whereas other choices such as hyperparameters changed only the scaling factor. Later works extended these ideas to joint model and dataset size scaling, such as those by Rosenfeld et al. (2019); Kaplan et al. (2020), and famously, the Chinchilla paper of Hoffmann et al. (2022) which introduced the notion of CO scaling measured using the loss on the training set. More recently, several efforts aim to predict downstream performance using scaling laws (Isik et al., 2025; Bhagia et al., 2024; Polo et al., 2025), but they focus on forecasting downstream metrics, whereas our work asks a fundamental question about how different skills exhibit different scaling behaviour.
@@ 6 Conclusion
Our analysis also opens intriguing directions for future study, such as understanding the relationship between compression and skills through the lens of COs, as well as understanding the flatness or curvature of IsoFLOP curves computed on the NLL compared to accuracy-based metrics.

### v3

@@ Results
In Figures 1(a) and 1(b), respectively, we show average results for the held-out code and knowledge QA datasets (in red). For comparison, we overlay the APE CO curve reported by Dubey et al. (2024) (in black). The figure shows a clear difference between knowledge QA and code: while knowledge QA prefers capacity compared to the APE curve – especially at lower compute scales – code tends to be more data-hungry. This pattern is present across all datasets belonging to the hypothesis split on which we first observed this difference, as well as for the held-out split. We show individual plots for all datasets in Figures 11, 12, 13 and 14.
To make the pattern more explicit across different scales, we consider the distribution of parameter count ‘residuals,’ or log-scale differences between the APE optima and the skill-optima for each skill. If at compute scale $B$, the residual is larger than 0, we say that the skill is capacity-hungry at $B$; if it is smaller than 0, we call it data-hungry at $B$. The resulting distributions, shown in Figure 1(c), confirm the pattern we observed before: the parameter count residuals for code are shifted substantially to the left with respect to the parameter count residuals for knowledge QA, indicating that the latter skill is more capacity-hungry. The same pattern holds for all datasets in the hypothesis and held-out splits. This strongly supports the claim that there is a difference in COs between knowledge QA and code, and shows that this observation is a generalisable phenomenon.
@@
In Figure 2, we show the results for the code and knowledge held-out datasets. Unsurprisingly, increasing the amount of skill-dependent data (lighter colors in the plot), improves performance for that specific skill. [^3] Increasing code data improved losses on code datasets beyond the next compute scale on the canonical datamix. This pattern is much clearer for code than for knowledge. We hypothesise that this is because code datapoints are easily identified and contain only code, whereas knowledge datasets are usually more dispersed as well as more difficult to label automatically. More relevant to our research questions is the fact that not only the losses, but also the COs shift. Again specifically for code, this pattern is clear in both the hypothesis and held-out sets – the more code data is included, the more capacity-hungry the CO becomes. For reference, in Appendix E, we provide individual plots for all code and knowledge QA datasets of all splits, across all ablations. This implies a fundamental relationship between the skill proportions in the datamix and COs – for a fixed compute budget, if the data requirement for a skill is satisfied by increasing its proportion in the datamix, we can afford to train a larger model on less data.
@@
However, we would like to assess whether or not the capacity- and data-hunger of knowledge and code are an artefact of the canonical datamix or if it is a more fundamental phenomenon. To do so, we use the same data to investigate the data-hunger of knowledge QA and code independently of the proportion of skill relevant data. To do so, we plot the CO parameter counts for each skill as a function of the proportion of skill relevant dta. In Figure 3, we show this for both knowledge QA and code in the same plot. [^4] The axis is aligned on proportions of skill-relevant data, not on the datamixes themselves. For knowledge QA datasets (purple lines), an x-value of 0.4 corresponds to a datamix with 40% knowledge, whereas for code datasets (green lines), it corresponds to results for the datamix with 40% code. From this figure it is clear that even correcting for the proportion of skill relevant data, knowledge is substantially more capacity-hungry than code, and increases its capacity-hunger more quickly as the proportion of skill relevant data increases. We hypothesise that this phenomenon, which demonstrates a fundamental difference between knowledge QA and code, is related to the fact that knowledge is harder to compress than code, requiring more capacity to memorise facts. The observation that COs for knowledge QA shift at a faster rate than code suggests a fundamental difference in their scaling, rather than an artefact of the datamix.
In line with this hypothesis, we observe a larger difference between knowledge and code when only Python code datasets are considered, as can be seen in Figure 4. We hypothesise that models at this scale may not be able to successfully compress ‘low-resource’ programming languages, though further research is required to confirm this.555 We also see in Figures 2 and 1 and Appendices E and I that datasets quantifying code have a lower loss value than those quantifying knowledge QA– across the board – despite large proportions of knowledge, which could suggest that code has lower unmodelable entropy compared to knowledge.
@@ 4.3.1 Aligning skill-COs
In Figure 5(b), we show the NLL for all validation sets, using the same axis as in Figure 5(a). The vertical dashed line indicates the optimal point from Figure 5(a). A first striking observation is that all the chosen validation sets track either knowledge QA or code quite closely: while the loss on FineWeb and FineWeb-Edu monotonically increases with the amount of code data, the opposite is true for The Stack. We do not see evidence of a mix of knowledge QA and code in the validation sets, which might have presented itself as a U-shaped curve, rather than the monotonically increasing or decreasing curves that we observe. To compute appropriate scaling laws, none of these validation sets would thus have resulted in COs aligned with both skills. This result implies that it is possible to align COs between skills, and that this should be done in practice to balance desired skills.
@@ 4.3.2 The impact of the validation set
Lastly, we study more explicitly how much variation in the predicted COs is to be expected depending on the choice of validation set. In Figure 6(a), considering The Stack validation set as ‘baseline,’ we can see that for some datamixes, the optimal parameter count is off by almost 50%. In Figure 6(c), we see that, for the smallest compute scale, for the canonical datamix there is over a 30% difference in optimal parameter count between the different validation sets. While the relative difference is not as stark at higher compute scales, it still exceeds 10% for the 3 largest compute scales. This experiment shows that choosing a validation set that adequately represents a mix of what the final model should capture – or perhaps even a validation set that more directly measures those skills, such as in our work – is critical to finding the right COs. While the difference is more pronounced at smaller scales, it persists even on the largest compute scale in our experiments.
On the other hand, Dubey et al. (2024) points out that IsoFLOP curves get flatter at high compute scales, so the tradeoff between model capacity and dataset size might not matter as much. In one of their other results, Dubey et al. (2024) uses a fitted sigmoid of the loss to estimate accuracies. If one were to take their analysis a step further by applying a sigmoid to an entire IsoFLOP curve to produce an IsoFLOP curve of accuracy rather than NLL, we hypothesise that this could make curves appear less flat if low compute scales are in the flat region of the sigmoid and high compute scales are in the linear region. We leave this intriguing question to future work.
@@ 5 Related work
The first large-scale empirical study of statistical scaling laws of neural networks was done by Hestness et al. (2017), who studied scaling laws for a variety of tasks using the LSTM Hochreiter and Schmidhuber (1997) and CNN Lecun et al. (1998) architectures. Their paper used power law fits to their experimental data, and they found that the only way to materially change the power law exponent was to change the task, whereas other choices such as hyperparameters changed only the scaling factor. Later works extended these ideas to joint model and dataset size scaling, such as those by Rosenfeld et al. (2019); Kaplan et al. (2020), and famously, the Chinchilla paper of Hoffmann et al. (2022) which introduced the notion of CO scaling measured using the loss on the training set. More recently, several efforts aim to predict downstream performance using scaling laws (Isik et al., 2025; Bhagia et al., 2024), but they focus on forecasting downstream metrics, whereas our work asks a fundamental question about how different skills exhibit different scaling behaviour. Related to these efforts is a concurrent work that characterises LLM skills as latent variables of the scaling model fit to benchmark scores of existing pretrained models, but differs from our work in its treatment of skills and crucially, that it does not study compute optimality Polo et al. (2025).
@@ 6 Conclusion
Our analysis also opens intriguing directions for future study, such as understanding the relationship between compression and skills through the lens of COs, as well as understanding the flatness or curvature of IsoFLOP curves computed on the NLL compared to accuracy-based metrics. Crucially, we note that the bulk of our analysis can be easily extended to additional skills using open source scaling suites, as it only requires running evaluations on pretrained models.